### v1

# CollabLLM: From Passive Responders to Active Collaborators

Abstract
Large Language Models are typically trained with next-turn rewards, limiting their ability to optimize for long-term interaction. As a result, they often respond passively to ambiguous or open-ended user requests, failing to help users reach their ultimate intents and leading to inefficient conversations. To address these limitations, we introduce CollabLLM, a novel and general training framework that enhances multiturn human-LLM collaboration. Its key innovation is a collaborative simulation that estimates the long-term contribution of responses using Multiturn-aware Rewards. By reinforcement fine-tuning these rewards, CollabLLM goes beyond responding to user requests, and actively uncovers user intent and offers insightful suggestions—a key step towards more human-centered AI. We also devise a multiturn interaction benchmark with three challenging tasks such as document creation. CollabLLM significantly outperforms our baselines with averages of 18.5% higher task performance and 46.3% improved interactivity by LLM judges. Finally, we conduct a large user study with 201 judges, where CollabLLM increases user satisfaction by 17.6% and reduces user spent time by 10.4%.
1 Introduction
Modern Large Language Models (LLMs) excel at generating high-quality single-turn responses when given well-specified inputs. However, real-world users often do not fully articulate their intents and sometimes initiate conversations with an imprecise understanding of their own needs (Taylor, 1968). As a result, users routinely refine their requests post hoc through iterative corrections, which can increase frustration, hinder effective task completion, and reduce conversational efficiency (Amershi et al., 2019; Zamfirescu-Pereira et al., 2023; Wang et al., 2024; Kim et al., 2024). Therefore, an open problem is to train models that actively guide users in clarifying and refining their intents, and helps them achieve their goals. This key challenge would improve user satisfaction and efficiency and streamline human-LLM interactions—especially as LLMs are being applied to real-world tasks that are increasingly complex and open-ended.
A notable limitation of established fine-tuning techniques, such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), is that they primarily reward LLMs for immediate, single-turn responses, reducing their incentive to seek clarification or assist users in refining their intents or preferences. As a result, commonly used LLMs tend to prioritize direct answers, even though seeking additional context would enhance task completion and increase user satisfaction (Kim et al., 2024).
Here we introduce CollabLLM, a novel and general training framework that improves the ability of LLMs to effectively collaborate with humans in multiturn scenarios (Gao et al., 2019; Balog & Zhai, 2023; Rahmani et al., 2023). The key innovation of CollabLLM is to promote LLMs’ forward-looking behavior that leads to long-term collaboration gains (Figure 1). We introduce a collaborative simulation module that samples future conversations with users to estimate the long-term impact of model responses across multiple turns, a measure we term the Multiturn-aware Reward (MR). The MR function evaluates responses by incorporating both extrinsic metrics, such as task-specific success, and intrinsic metrics, such as efficiency, to holistically assess collaboration quality (cf. Section 3). By fine-tuning with established RL algorithms (Rafailov et al., 2023; Schulman et al., 2017) on MRs, CollabLLM promotes responses that lead to better task completion and efficiency in later conversation stages. As shown in Figure 2b, the fine-tuned model goes beyond simply responding to user requests in Figure 2a—it actively collaborates by asking follow-up questions about the writing tone, generating targeted content about the role of optimism, and offering insightful suggestions such as adding anecdotes.
We also introduce three challenging multiturn tasks for training and evaluation in simulated environments: MediumDocEdit-Chat, BigCodeBench-Chat, and MATH-Chat, which respectively encompass document creation, code generation, and multiturn question answering. On the three test sets, our approach improves task accuracy metrics by 18.5% and interactivity by 46.3% on average compared to our best baselines, according to LLM judges. Beyond the tasks that the CollabLLMs are fine-tuned on, we show CollabLLMs are highly generalizable to other data domains.
Moreover, we perform a large-scale and real-world user study with 201 Amazon Mechanical Turkers (MTurkers), who are asked to write documents with the help of anonymous AI assistants, either CollabLLM or non-collaboratively trained LLMs. CollabLLM achieves impressive improvement with 17.6% increase in user satisfaction and yield user time savings of 10.4% on average. The qualitative analysis from MTurkers confirms our observations: non-collaboratively-trained LLMs passively agree with users, while CollabLLM actively provide insightful questions and suggestions to guide writing processes.
2 Problem Formulation
In contrast to many existing tasks that are single-turn and require no human involvement beyond the initial query, our problem formulation reflects a real-world setting in which a user’s underlying (implicit) goal is defined as in a multiturn conversational task. The conversation unfolds over multiple turns , where is the user input and is the model’s response at each turn , where is the number of turns in the conversation.
At the -th turn, the model generates its response based on the previous conversation turns and the current user response . For simplicity, we define historical conversation at -th turn as , therefore, . The objective is to generate a sequence of model responses that effectively and efficiently achieve for goal , e.g., answering a math question, where goal achievement is assessed based on user satisfaction or an external evaluation function, such as accuracy by LLM judge. Formally, we define the objective as , where incorporate the achievement of task success and user experience factors such as time cost.
3 Unified Collaborative LLM Training
Key Motivations. Established LLM training frameworks, such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), focus on maximizing immediate rewards for single-turn tasks. This cause a misalignment between their single-turn objective and real-world multiturn objective . Precisely, the model’s accumulative single-turn reward may not imply a higher final reward . In fact, achieving high single-turn rewards at each turn may not imply a higher final reward. For example, consider a task where the user’s goal is to write an engaging article. A model trained with traditional RLHF might generate isolated responses, like drafting an introduction or listing conclusions. While these responses are helpful in isolation, they fail to consider how the sections flow together, resulting in an article that might not be cohesive and aligned with the user’s goal.
Instead, effective multiturn collaboration requires model responses that optimally contribute to the final reward. The model should aim to align its responses with the user’s goal by considering their impact on the entire conversation trajectory . In the previous example, instead of generating a conclusion, asking, “Should I maintain an engaging tone in the conclusion like the introduction?” offers better long-term alignment with the goal.
3.1 Multiturn-aware Rewards
In Figure 1, our key insight is that effective multiturn collaboration relies on forward-looking strategies. Given a context , the model should consider how its response influences the subsequent turns of the conversation. To capture this, we design a collaborative simulation module to estimate this impact. By fine-tuning to distinguish between potential future conversations resulting from different responses, the model generates responses that align better with the overarching goal .
This high-level design naturally aligns with causal effect estimation (Pearl, 2009; Pearl et al., 2016), which evaluates the interventional effects of an action in sequential decision-making. Appendix A provides further discussion on the connection between causal effect estimation and our approach. More specifically, we define the Multiturn-aware Reward:
Multiturn-aware Reward (MR): The multiturn-aware reward for model response at the -th turn is given by:
where denotes the conversation history up to and including the -th turn, and represents the forward trajectory of turns following the -th turn. The distribution models the possible forward conversations conditioned on the prior conversation history.
However, computing Equation 1 remains challenging as it requires the following components: (a) A conversation-level reward function, , for evaluating an arbitrary multiturn conversation , and (b) a sampling strategy for obtaining forward conversations , which represents the forward conversation distribution. We elaborate on the two components in Section 3.1.1 and 3.1.2.
3.1.1 Conversation-level Reward Function
We approximate the conversation-level reward with a combination of extrinsic (goal-specific) and intrinsic (goal-agnostic) metrics:
where focuses on task-specific success, and evaluates user experience including efficiency and engagement.
Extrinsic Reward measures how well the conversation achieves the user’s goal . Formally:
(3) where extracts the final solution or response from the conversation , especially for tasks requiring revisions or multi-step answers. is the reference solution for the goal , e.g., the ground truth solution for a math problem. And evaluates task-specific metrics like accuracy or similarity. This ensures the conversation contributes directly to achieving the desired goal.
Intrinsic Reward prioritizes conversations that enhance user experience, defined as:
(4) where we encourage conversational efficiency by penalizing excessive tokens that users read and write, with controlling the penalty severity. This efficiency measure is bounded by 1 to maintain balance with other metrics. The second term, , is assigned by an LLM-based judge (Zheng et al., 2023) on a 0–1 scale, evaluating user-valued objectives such as engagement / interactivity. Notably, additional conversational aspects, such as clarity, can be further integrated into the objective.
The conversation-level reward incorporates task-specific and human-centered metrics, encouraging the model to balance goal achievement, efficiency, and engagement.
3.1.2 Forward Sampling
To compute Eq. 1, we require samples from , the distribution of forward conversation conditioned on the conversation history. A simple approach is to use Monte Carlo sampling, where the conversation is extended turn-by-turn until it concludes. However, this can be computationally expensive for computing reward for every model response. For a scalable approximation, we introduce a window size as a hyperparameter to limit the maximum number of forward turns considered in . This reduces the computational cost while maintaining sufficient context.
More importantly, while real-world conversations could be gathered from human participants, sampling multiple forward conversations during training is costly and impractical. To further reduce cost and ensure scalability, we introduce a user simulator .
User Simulator:. A user simulator is a function that maps a given conversation history to a user response . Specifically, generates a probabilistic distribution over possible user responses conditioned on the conversation history , simulating realistic user behavior.
Specifically, we prompt an LLM to role-play as users, explicitly asking the LLM to follow the same language style as the previous user turns, and injecting typical user behaviors. The user simulator operates with an implicit goal , which it seeks to achieve over the course of the conversation. This design emulates real-world scenarios where users may have evolving needs, limited background knowledge, or require clarification, resulting in naturally unfolding multiturn conversations (Park et al., 2024).
3.2 Optimization & Synthetic Datasets
With the conversation-level reward function and forward sampling strategy, we can compute MR for any model response without requiring an additional reward model, which is often costly and slow to train. Unlike traditional single-turn reward approaches, MR explicitly accounts for the impact of a response on future conversations, promoting long-term collaboration.
Further, we employ reinforcement learning (RL) methods such as PPO (Schulman et al., 2017) and DPO (Rafailov et al., 2023) to guide the model in navigating complex conversations. By optimizing for higher MR, the model learns to generate responses that enhance overall effectiveness and efficiency by the end of the conversation.
Moreover, MR can generate high-quality synthetic conversations (cf. Figure 8 in Appendix B) for both supervised fine-tuning (SFT) and DPO. For SFT, it iteratively selects top-ranked responses to build realistic, goal-directed conversation histories. For DPO, it constructs pairwise comparisons by ranking responses at each turn, distinguishing “chosen” and “rejected” pairs based on MR scores. The generated synthetic data aligns with multiturn objectives.
Overall, CollabLLM enables scalable dataset generation and online RL training without human annotation, making it generalizable across diverse tasks. In Appendix 7, we compare CollabLLM with related prompting- and training-based approaches, highlighting its contributions.
4 Experimental Setup***Dataset and training details in Appendix B; all prompts (e.g., prompts of user simulator and LLM judges) in Appendix C.
For fine-tuning and evaluation, we create three multiturn datasets using publicly available data across diverse domains (Hendrycks et al., 2021; Zhuo et al., 2024; Chiusano, 2024): collaborative document editing, coding problem assistance, and multiturn mathematics problem solving.
To build a multiturn environment (Figure 3), we employ GPT-4o-mini as a user simulator LLM to role-play realistic user behaviors, given the target problem and conversation history. Our simulation-based evaluations are designed to closely mimic real-world interactions (Park et al., 2024). Unlike traditional single-turn tasks, our setup requires dynamic interactions over multiple turns to achieving a goal. The three interactive datasets are:
MediumDocEdit-Chat: Document editing requires iterative feedback and refinements across multiple turns to ensure coherence and alignment with user intent. We sample 100 Medium articles as goal documents, which are summarized into target problems to guide the user simulator. After each interaction, task performance is evaluated using the BLEU score, measuring similarity between the extracted document and the original articles.
BigCodeBench-Chat: Coding tasks inherently require multiturn interactions, such as clarifying requirements and debugging. We sample 600 coding problems from BigCodeBench (Zhuo et al., 2024) as the target problems given to the user simulator. For evaluation, we compute the average Pass Rate (PR) of code at the end of the interactions.
MATH-Chat: Math problem solving often requires addressing implicit assumptions, verifying intermediate steps, and clarifying reasoning. We sample 200 level-5 math problems from MATH (Hendrycks et al., 2021) to prompt the user simulator, which interacts with the LLMs. Task success is measured by the accuracy (ACC) of the final solution, as evaluated by an LLM judge.
In addition to the above task-specific metrics, we incorporate two task-agnostic scores across all datasets: 1) Average Token Count, which quantifies the average number of tokens generated by the LLM per conversation, reflecting interaction efficiency. 2) Interactivity (ITR), which evaluates engagement levels using an LLM judge (Claude-3.5-Sonnet), with scores rescaled to the range [0, 1].
| MediumDocEdit-Chat | BigCodeBench-Chat | MATH-Chat | |||||||
| BLEU | #Tokens | ITR | PR | #Tokens | ITR | ACC | #Tokens | ITR | |
| Base | 32.2 | 2.49 | 46.0 | 9.3 | 1.59 | 22.0 | 11.0 | 3.40 | 44.0 |
| Proactive Base | 35.0 | 2.18 | 62.0 | 11.0 | 1.51 | 33.7 | 12.5 | 2.90 | 46.0 |
| SFT | 35.2 | 2.21 | 68.0 | 11.7 | 1.35 | 42.0 | 13.5 | 2.88 | 58.0 |
| PPO | 38.5 | 2.00 | 78.0 | 14.0 | 1.35 | 40.7 | 13.0 | 2.59 | 52.0 |
| Offline DPO | 36.4 | 2.15 | 82.0 | 12.3 | 1.35 | 46.7 | 15.5 | 2.40 | 50.0 |
| Online DPO | 36.8 | 2.00 | 92.0 | 13.0 | 1.31 | 52.0 | 16.5 | 2.37 | 60.0 |
| Rel. Improv. | 5.14% | 8.25% | 48.3% | 18.2% | 13.2% | 54.3% | 32.0% | 18.3% | 36.4% |
Fine-tuning CollabLLMs. CollabLLMs are based on Llama-3.1-8B (Llama Team, 2024) with LoRA finetuning (Hu et al., 2022). We train four model variants: 1) Offline models: SFT and Offline DPO are fine-tuned on pre-generated multiturn conversational datasets guided by Multiturn-aware Rewards (MR) (cf. Section 3.2). 2) Online models: PPO and Online DPO are further trained from the SFT and Offline DPO models, respectively. The model during online fine-tuning is involved in the collaborative simulation to compute MRs, which, in turn, dynamically adjust the model preference.
Baselines. We compare CollabLLMs against (1) the pretrained Llama-3.1-8B (Base), (2) the base model with proactive prompt engineering (Proactive Base), which encourages follow-up and clarification questions.
| Action-level Accuracy | Macro Metric | |||
| Ambiguous | Non-Ambiguous | Accuracy | F1 | |
| GPT-4o | 15.44% | 95.60% | 55.52% | 56.62% |
| Base (Llama-3.1-8B) | 16.26% | 90.40% | 53.33% | 53.31% |
| CollabLLM | 52.84% | 72.32% | 62.58% | 55.08% |
5 Results of Simulated Experiments
We present the results in Table 1 and the takeaways are:
Prompt engineering is helpful, but limited in terms of performance gains and flexibility. Proactive Base improves base model performance by encouraging follow-up questions and clarifications. For example, it increases BLEU on MediumDocEdit-Chat from 32.2% to 35.0% and reduces read tokens by 0.31k compared to the base model. However, these gains are modest and do not fully address the challenges of multiturn collaboration. We observe that prompting strategies remain rigid, relying on predefined instructions rather than adapting dynamically to user needs. For instance, the model sometimes asks clarification questions even when unnecessary, leading to redundant interactions that disrupt conversation flow.
CollabLLM increases task performance, conversational efficiency, and engagement. CollabLLM achieves 18.5% superior task-specific performance, 13.3% more efficient conversations, and 46.3% enhanced interactivity compared to the best baselines. We highlight that CollabLLM engage in more meaningful collaborations, with ITR shows substantial gains. For MediumDocEdit-Chat, the Online DPO model increases ITR from 0.46 to 0.92. Moreover, our framework significantly improves conversational efficiency by minimizing the content users need to review to arrive at the final solution. For MATH-Chat, Online DPO decreases token count per conversation by 1.03k compared to the base model.
5.1 Ablations on Reward Mechanisms (Figure 4)
To investigate how components contribute to CollabLLM’s superior performance, we conduct an ablation study focusing on the reward mechanisms used during fine-tuning. We evaluate the following reward mechanisms:
Variants of Multiturn-aware Reward: We vary the forward sampling window size to assess their ability to capture long-term conversational effects through simulated collaborations.
Immediate Rewards evaluate the model’s immediate response based on: 1) Helpfulness: Assessed by an LLM judge; 2) Extrinsic Reward: Focuses on task-specific metrics like BLEU while ignoring intrinsic factors such as efficiency; 3) Extrinsic + Intrinsic Reward: Combines task-specific metrics with efficiency and interactivity measures. This can be seen as a special case of the multiturn-aware reward function with .
We present results in Figure 4. Interestingly, expanding the forward sampling window within the range generally enhances performance and efficiency by better capturing future interactions. Notably, MR with balances the gains and additional costs to conduct forward sampling, making it well-suited for large-scale fine-tuning. In contrast, immediate rewards, even with extrinsic and intrinsic components, fall short as they ignore long-term impact. These findings validate the positive impact of the forward sampling strategy in MRs.
5.2 Case Study (Figure 5 & 6)
We now offer a deeper insight into CollabLLM’s behavior as shown in Figure 5. In this example, the user request to tokenize a text file is inherently open-ended due to unspecified factors, such as the NLTK environment, tokenizer selection, and optional preprocessing steps. The base LLM makes several arbitrary assumptions, applying lowercase conversion and stopword removal without user confirmation. The user simulator later corrects these assumptions, but the final solution remains incorrect due to missing stopwords. In contrast, CollabLLM actively clarifies user intent by seeking confirmation on key decisions, ensuring an aligned final solution with a 100% Pass Rate. This approach also reduces user effort with lower token usage.
In Figure 6, we compare different reward mechanisms for responses A and B of Figure 5, to confirm that these rewards work as intended. The helpfulness rewards favor response A due to its seemingly more well-round output. Extrinsic rewards assign zero scores to both, as A provides an incorrect solution and B defers answering. Extrinsic + Intrinsic rewards slightly favor B for efficiency and engagement. Interestingly, MR assigns significantly higher rewards to B, especially at and , since the response obtains useful information and provide a precise answer within the future interaction window.
| Model | Strengths | Weaknesses |
| Base | “Follows great instruction and does exactly what I’m asking it to do.”, “It can create a nice form of an outline to work with.” | “The AI just agreed with me on pretty much everything. There was no discussion”, “I didn’t really like that it kept coming up with different options” |
| Proactive Base | “It is very organized and it actually asks you for feedback after writing the revision.” | “The AI seemed to be very redundant and asked me the same questions over and over.” |
| Collab LLM | “Asking questions and making you think of things you never thought of”, “The AI really helped me with focusing on one part of the story at a time.”, “It helped really well to navigate what to say and what information is needed” | “The AI assistant was not up to date enough to help with this recent sporting event. The AI assistant also asked me to repeat information I had already given it.” |
5.3 Model Generalization (Table 2)
Modern foundation models are expected to generalize across a diverse range of tasks beyond their training domain. A key question is whether collaborative behaviors learned by CollabLLM during fine-tuning transfer effectively to new tasks without additional adaptation.
We assess CollabLLM, trained with online DPO on BigCodeBench-Chat (the coding assistance task), on Abg-CoQA (Guo et al., 2021), a question-answering (QA) benchmark where questions are labeled as ambiguous or non-ambiguous (cf. Appendix D). We categorize the model’s responses into two actions—asking a clarifying question or providing a direct answer—and evaluate action-level accuracy within each question type. As shown in Table 2, GPT-4o and Llama-3.1-8B rarely ask clarifying questions regardless of ambiguity. In contrast, CollabLLM proactively asks questions about 50% of the time while maintaining high accuracy on unambiguous inputs. This behavior leads to the highest Macro Accuracy across both ambiguous and non-ambiguous sets and improves Macro F1 over the base model, while leaving room for further improvement against GPT-4o. These results suggest that CollabLLM effectively generalizes its learned collaborative strategies beyond its training domain.
6 Real-world User Study
Setup. We conduct a large-scale user study using Amazon Mechanical Turk with 201 participants. Each participant is assigned a document type—randomly selected to be either blog post, creative writing, or personal statement—and chooses a topic from a predefined set. To simulate real-world scenarios where users have only a rough idea of the task, they are first asked to provide brief responses to topic-related questions. Participants then engage in at least eight turns of conversation with an anonymized AI assistant, which can be Base, Proactive Base, or CollabLLM. Every three turns, they provide an interaction rating based on their experience so far. After the conversation, participants rate the final document quality and overall interaction. All ratings are in a scale from 1 to 10. We also record the total interaction duration to assess efficiency. The detailed user study setup is provided in Appendix E.
Quantitative Results (Figure 7). Across multiple metrics, CollabLLM consistently outperforms the baselines. It achieves an average document quality score of 8.50. Specifically, 91.4% of participants rate CollabLLM’s document quality as “good” (score 8–9), and 56.9% as “very good” (score 9–10), compared to 88.5% and 39.3% for Base (Llama-3.1-8B), respectively. Similarly, 63.8% of participants find CollabLLM highly engaging, while only 42.6% report the same for Llama-3.1-8B.
Interestingly, for multiturn interaction, the Base model shows a declining trend in ratings from turns 6–9, indicating reduced user experience in longer conversations. In contrast, both CollabLLM and Proactive Base exhibit increasing ratings over time, with CollabLLM consistently achieving higher average ratings every three turns compared to Proactive Base. This suggests that CollabLLM maintains sustained engagement more effectively.
Moreover, CollabLLM improves task efficiency, reducing time spent by 10.4% compared to the Base model and by 15.6% relative to Proactive Base. While Proactive Base is prompted to maintain conciseness, it frequently asks unnecessary questions, causing lower efficiency. In contrast, CollabLLM strikes a more streamlined user experience.
Qualitative Results (Table 3). We collected a total of 180 strengths and 180 weaknesses across the three models. Table 3 presents representative feedback, while we summarize here the mddels’ strengths and weaknesses: The base model generates coherent content while effectively follow user instructions, but it sometimes struggles with maintaining context in long texts, and can be overly verbose or repetitive in its responses. Proactive Base excels in responsiveness and adapting to user input but struggles with memory retention, and could produce repetitive or overly structured content. On the other hand, CollabLLM is highly engaging, effectively guiding users through writing, adapting seamlessly to feedback. However, users also point out that CollabLLM can occasionally feel bland, lack of up to date information, and require additional effort to personalize the output. Overall, CollabLLM enhances collaboration by guiding users through an interactive and iterative refinement process, yet future improvements should focus on increasing personalization, creativity, and real-time knowledge integration to further optimize human-LLM collaboration.
7 Related Work
Non-collaborative LLM training. Existing LLM training frameworks, including pre-training, supervised fine-tuning (SFT), and reinforcement learning (RL) (Rafailov et al., 2023; Schulman et al., 2017; Ouyang et al., 2022; Lee et al., 2024), primarily optimize for next-turn response quality. Standard RL methods such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) apply rewards to individual model responses without accounting for their long-term impact on conversation trajectories. While effective for single-turn objectives, these approaches fail to capture how responses influence user intent discovery and long-term task success (Amershi et al., 2019; Zamfirescu-Pereira et al., 2023; Wang et al., 2024; Kim et al., 2024; Lu et al., 2024).
Prompting techniques for multiturn interaction. Prior work has explored prompting strategies to enhance LLM interactivity, particularly for clarification questions (Keh et al., 2024; Mu et al., 2023; Zhang & Choi, 2023; Chi et al., 2024; Kim et al., 2023; Deng et al., 2023b; Zhao & Dou, 2024) and mixed-initiative dialogues (Deng et al., 2023a; Chen et al., 2023; Liao et al., 2023). For instance, Mu et al. (2023) prompt LLMs to ask clarification questions when code generation requests are ambiguous. However, such prompting-based approaches are constrained by predefined interaction patterns, limiting adaptability across different tasks and conversation stages. Moreover, their reliance on fixed prompts reduces generalization, as demonstrated in our experiments where proactive prompting fails to match the effectiveness of our fine-tuned models.
Learning-based methods for multiturn interaction.
LLMs for generating clarification questions: Beyond prompting, prior studies have explored supervised fine-tuning (Andukuri et al., 2024), RL fine-tuning (Chen et al., 2024; Zamani et al., 2020; Erbacher & Soulier, 2023), and active learning (Pang et al., 2024) to train models to ask clarification questions. For example, Chen et al. (2024) use Direct Preference Optimization (DPO) to encourage models to identify ambiguous user queries and request clarifications. However, like prompting approaches, these methods primarily focus on clarification questions and do not generalize to broader multiturn collaboration strategies.
Multiturn training for LLMs: Several studies extend RLHF to multiturn settings by optimizing trajectory-level rewards (Shani et al., 2024; Zhou et al., 2024; Gao et al., 2024; Shi et al., 2024b). Other works (Xu et al., 2023; Deng et al., 2024) leverage self-chat or self-play to enhance model adaptation to dialogue. See Zhang et al. (2025) for a comprehensive survey. However, these methods primarily rely on post-hoc trajectory-level data, learning from observed conversations rather than explicitly modeling the causal effect of individual responses on long-term task success (see Appendix A for further details). Additionally, they often overlook open-ended tasks such as document generation (Faltings et al., 2023; Jiang et al., 2024), where real-world user responses can be highly variable, and users may have limited capacity to read and refine lengthy model outputs.
In Table 4, we compare CollabLLM with related methods across four key dimensions. CollabLLM is a general, user-centric, and multiturn-aware framework that leverages more accurate reward estimation to better align with real-world objectives, enhancing user satisfaction and streamlining human-LLM interactions.
8 Conclusion
Multiturn human-LLM collaborations are increasingly prevalent in real-world applications. We believe future foundation models should act as collaborators rather than passive responders, actively uncovering user intents in open-ended and complex tasks—an area where current LLMs fall short. The key insight of CollabLLM is making LLMs more multiturn-aware by using forward sampling to estimate the long-term impact of responses. Through extensive simulated and real-world evaluations, we demonstrate that CollabLLM is highly effective, efficient, and engaging, while also generalizing well to new tasks and interactions, advancing the frontiers of human-centered LLMs.
Impact Statement
This paper presents work aimed at making AI more user- and human-centric, which, in our view, yields a positive societal impact. Most current work on AI and its evaluation focuses on fully automated tasks, with no user involvement in solving the task or optimization for a collaborative experience with users. This has serious societal drawbacks, given issues such as AI hallucinations (Huang et al., 2025), biases (Gallegos et al., 2024), and unsafe language (Shi et al., 2024a) that arise from a lack of human oversight. The common focus on having AI models autonomously complete tasks also ignores the reality that many scenarios have humans present regardless of the level of automation, and that not priming AI models to proactively seek human help, feedback, or clarifications misses an opportunity to make generative AI more accurate, effective, and safe. This consideration would also help increase the adoption of AI in safety-critical scenarios, such as medical decision-making tasks (Liu et al., 2024), in which we believe AI models should be inclined to seek confirmation or verification (Gero et al., 2023) from an expert in case of uncertainty—a behavior that is mostly absent in current state-of-the-art LLMs.
This work presents one of the first attempts to train LLMs in such human-centric environments. To promote future research in this societally beneficial direction, we plan to release all the code, models, data, benchmarks, and user simulators described in this work.
The data collected in our study involves human participants recruited through Mechanical Turk. We took several measures to ensure the privacy of these workers in the document creation tasks. First, we asked workers to confirm that they were willing to share the text they wrote as part of a public dataset. Second, we urged them not to include any personally identifiable information (PII) in their writings and to focus only on topics of public knowledge or fictitious stories. Third, we scanned the collected data to ensure that no PII was included. For the final version of the dataset, we will recruit additional workers to manually review each collected conversation to ensure that no PII or other safety issues (e.g., offensive language) exist in the data. Mechanical Turk workers were paid $10 per conversation. Given that conversations averaged 28.4 minutes, including break times, this means workers were paid more than $20 per hour on average—above the minimum wage in the country where the data was collected.
Acknowledgments
We thank Vishal Chowdhary, Hoifung Poon, Swadheen Shukla, Chandan Singh and Chenglong Wang, as well as members of the Deep Learning and Health Futures groups at Microsoft Research for helpful discussions. We thank lab members in Leskovec and Zou’s labs for discussions and for providing feedback.
Appendix A Supplementary Discussion
A.1 Connection Between Multiturn-aware Reward and Causal Inference
Our approach naturally aligns with causal inference principles, as it aims to quantify how a model’s response influences the future trajectory of a conversation. This aligns with the fundamental goal of causal effect estimation, which seeks to isolate the impact of an intervention—in this case, a model response—on long-term outcomes.
From a causal perspective, given a conversation history at turn , the causal effect of a model response on the final conversation trajectory can be expressed using front-door adjustment (Pearl, 2009; Pearl et al., 2016):
This equation captures the expected long-term reward of a conversation conditioned on the model’s response at turn . It explicitly accounts for how intervenes in the conversation, influencing future turns and, ultimately, task success.
A.2 Distinction from Other Multiturn Training Frameworks
Existing multiturn trajectory-based training frameworks (Shani et al., 2024; Zhou et al., 2024; Gao et al., 2024) primarily rely on learning from observed trajectory-level rewards. These methods estimate the utility of responses by assigning rewards post hoc to completed conversations, training models to imitate historically successful responses. However, this approach is fundamentally correlational—it captures statistical associations between responses and final outcomes but does not explicitly model how different responses causally influence future turns.
In contrast, our Multiturn-aware Reward (MR) framework intervenes on model responses and uses forward sampling to explicitly simulate alternative future trajectories. This enables the model to estimate the counterfactual impact of different responses and adjust its behavior accordingly. By leveraging causal effect estimation, MR training moves beyond passive imitation of high-reward conversations and instead actively optimizes responses to maximize long-term goal attainment. This distinction is crucial in dynamic human-LLM interactions where user needs evolve throughout a conversation.
Appendix B Dataset and Training Details
B.1 Dataset Generation for Offline Training
The Multiturn-aware Reward (MR) function enables the generation of high-quality synthetic conversation datasets for training. Given a user query, multiple LLM responses are sampled and ranked based on their MR scores, with higher-ranked responses designated as Chosen and lower-ranked as Rejected. To simulate natural conversational flow, the first turn from the chosen response’s forward interaction window is appended to the prompt for the next turn, iteratively extending the conversation until completion. Solid red arrows denote data collection for Supervised Fine-Tuning (SFT), while dashed blue arrows indicate preference data construction for Direct Preference Optimization (DPO). This approach systematically curates multiturn conversations that enhance both response quality and collaborative efficiency, both of which are explicitly captured by MR.
Given (1) a user simulator LLM, e.g., GPT-4o-mini, (2) an assistant LLM, GPT-4o, and (3) arbitrary tasks with defined task-specific metric, we can simulated and generate high-quality conversations following Figure 8. We create the following training datasets in this simulated environments.
| # Train | # Turns | Average # Turns | Chosen MR | Rejected MR | |
| MediumDocEdit-Chat | 500 | 2,303 | 4.61 | 0.312 0.104 | 0.246 0.113 |
| BigCodeBench-Chat | 500 | 2,627 | 5.25 | 0.494 0.621 | 0.207 0.763 |
| MATH-Chat | 500 | 2,527 | 5.05 | 0.863 0.524 | 0.547 0.502 |
B.2 Training Details
We provide the hyperparameters for CollabLLM fine-tuning in Table 6.
Notably, CollabLLM relies on a minimal set of hyperparameters, using the same window size and sample size for computing MRs across multiple datasets. The penalty factor on token count, , is set lower for MediumDocEdit-Chat compared to BigCodeBench-Chat and MATH-Chat, as document lengths in MediumDocEdit-Chat can vary significantly and may be easily bounded by 1 in Eq. 4 if is too large.
| LoRA Configuration | |
| Rank | 32 |
| Scaling factor | 16 |
| Dropout | 0.1 |
| Bias | False |
| Fine-Tuning Hyperparameters | ||||
| SFT | Offline DPO | Online DPO | PPO | |
| Learning rate | 1e-5 | 5e-6 | 5e-6 | 2e-6 |
| Total batch size | 64 | 64 | 32 | 64 |
| Number of epochs | 3 | 8 | 1 | 5 |
| CollabLLM-specific Hyperparameters | |||
| MediumDocEdit-Chat | BigCodeBench-Chat | MATH-Chat | |
| Window size | 2 | 2 | 2 |
| Sample size for MR | 3 | 3 | 3 |
| Penalty | 1e-4 | 5e-4 | 5e-4 |
Appendix C Prompts
C.1 User Simulator
C.2 Prompt for Proactive Base
C.3 System Prompt
C.4 Interactivity Metric by LLM Judge
C.5 Helpfulness Reward by LLM Judge
Appendix D Question Template and Example on Abg-CoQA
We use the following prompt format for the LLMs to answer the question given a story.
For example:
The label of the above question is ambiguous since the user’s query about ‘‘Where did they go when they left?’’ could mean ‘‘Where did they go when they left the zoo?’’ or ‘‘Where did the grandparents go when they left me?’’.
Appendix E User Study
We provide screenshots of the interface used for human participants to interact with the AI assistants. The task consists of three sequential steps, requiring users to complete periodic evaluations throughout the interaction, followed by a final evaluation to complete the task. All data collection is fully anonymized to ensure user privacy.

### v2

@@
 Section 3).
-By fine-tuning with established RL algorithms (Rafailov et al., 2023; Schulman et al., 2017) on MRs, CollabLLM promotes responses that lead to better task completion and efficiency in later conversation stages.
+By fine-tuning with RL algorithms (Rafailov et al., 2023; Schulman et al., 2017) on MRs, CollabLLM promotes responses that lead to better task completion and efficiency in later conversation stages.
 As shown in Figure 2b, the fine-tuned model goes beyond simply responding to user requests in Figure 2a—it actively collaborates by asking follow-up questions about the writing tone, generating targeted content about the role of optimism, and offering insightful suggestions such as adding anecdotes.
@@
 We approximate the conversation-level reward with a combination of extrinsic (goal-specific) and intrinsic (goal-agnostic) metrics:
-where focuses on task-specific success, and evaluates user experience including efficiency and engagement.
+where focuses on task success, and evaluates user experience including efficiency and engagement.
 Extrinsic Reward measures how well the conversation achieves the user’s goal . Formally:
@@
 In Appendix 7, we compare CollabLLM with related prompting- and training-based approaches, highlighting its contributions.
-4 Experimental Setup***Dataset and training details in Appendix B; all prompts (e.g., prompts of user simulator and LLM judges) in Appendix C.
+4 Experimental Setup***Dataset and training details in Appendix B; all prompts (e.g., prompts of user simulator and LLM judges) in Appendix D.
 For fine-tuning and evaluation, we create three multiturn datasets using publicly available data across diverse domains (Hendrycks et al., 2021; Zhuo et al., 2024; Chiusano, 2024): collaborative document editing, coding problem assistance, and multiturn mathematics problem solving.
@@
 In addition to the above task-specific metrics, we incorporate two task-agnostic scores across all datasets: 1) Average Token Count, which quantifies the average number of tokens generated by the LLM per conversation, reflecting interaction efficiency.
-2) Interactivity (ITR), which evaluates engagement levels using an LLM judge (Claude-3.5-Sonnet), with scores rescaled to the range [0, 1].
+2) Interactivity (ITR), which evaluates engagement levels using an LLM judge (Claude-3.5-Sonnet), with scores rescaled to an upper bound of 1.
 | MediumDocEdit-Chat | BigCodeBench-Chat | MATH-Chat | |||||||
@@
 For instance, the model sometimes asks clarification questions even when unnecessary, leading to redundant interactions that disrupt conversation flow.
-CollabLLM increases task performance, conversational efficiency, and engagement.
+CollabLLM improves task performance, efficiency, and engagement.
 CollabLLM achieves 18.5% superior task-specific performance, 13.3% more efficient conversations, and 46.3% enhanced interactivity compared to the best baselines.
@@
 We assess CollabLLM, trained with online DPO on BigCodeBench-Chat (the coding assistance task), on Abg-CoQA (Guo et al., 2021), a question-answering (QA) benchmark where questions are labeled as ambiguous or non-ambiguous (cf.
-Appendix D).
+Appendix E).
 We categorize the model’s responses into two actions—asking a clarifying question or providing a direct answer—and evaluate action-level accuracy within each question type.
@@
 We also record the total interaction duration to assess efficiency.
-The detailed user study setup is provided in Appendix E.
+The detailed user study setup is provided in Appendix F.
 Quantitative Results (Figure 7).
@@
 Standard RL methods such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) apply rewards to individual model responses without accounting for their long-term impact on conversation trajectories.
-While effective for single-turn objectives, these approaches fail to capture how responses influence user intent discovery and long-term task success (Amershi et al., 2019; Zamfirescu-Pereira et al., 2023; Wang et al., 2024; Kim et al., 2024; Lu et al., 2024).
+While effective for single-turn objectives, these approaches fail to capture how responses influence user intent discovery and long-term task success (Amershi et al., 2019; Zamfirescu-Pereira et al., 2023; Wang et al., 2024; Kim et al., 2024).
 Prompting techniques for multiturn interaction.
@@
 For example, Chen et al.
-(2024) use Direct Preference Optimization (DPO) to encourage models to identify ambiguous user queries and request clarifications.
+(2024) use Direct Preference Optimization (DPO) to encourage models to request clarifications.
 However, like prompting approaches, these methods primarily focus on clarification questions and do not generalize to broader multiturn collaboration strategies.
-Multiturn training for LLMs: Several studies extend RLHF to multiturn settings by optimizing trajectory-level rewards (Shani et al., 2024; Zhou et al., 2024; Gao et al., 2024; Shi et al., 2024b).
-Other works (Xu et al., 2023; Deng et al., 2024) leverage self-chat or self-play to enhance model adaptation to dialogue.
-See Zhang et al.
-(2025) for a comprehensive survey.
-However, these methods primarily rely on post-hoc trajectory-level data, learning from observed conversations rather than explicitly modeling the causal effect of individual responses on long-term task success (see Appendix A for further details).
-Additionally, they often overlook open-ended tasks such as document generation (Faltings et al., 2023; Jiang et al., 2024), where real-world user responses can be highly variable, and users may have limited capacity to read and refine lengthy model outputs.
+Multiturn training for LLMs: Recent benchmarks (Abdulhai et al., 2023; Kwan et al., 2024) evaluate LLMs’ performance in multiturn settings, measuring the goal orientation and planning capabilities of models across interactions.
+Several studies extend RLHF to multiturn settings by optimizing trajectory-level rewards (Shani et al., 2024; Zhou et al., 2024; Gao et al., 2024; Shi et al., 2024b; Zhang et al., 2025).
+Other works (Xu et al., 2023; Deng et al., 2024) leverage self-chat or self-play to enhance model adaptation.
+However, these methods primarily rely on post-hoc trajectory-level data, learning from observed conversations rather than explicitly modeling the causal effect of individual responses on task success (see Appendix A for further explanations).
+Additionally, they often overlook open-ended tasks such as document generation (Faltings et al., 2023; Jiang et al., 2024), where user responses can be highly diverse, and users may have limited capacity to read and refine lengthy model outputs.
+User simulators for enhancing AI systems.
+Recent works employ user simulators to enhance dialogue systems (Shi et al., 2019; Tseng et al., 2021) and LLMs (Hong et al., 2023; Hu et al., 2023; Faltings et al., 2023).
+Recently, Hong et al.
+(2023) leverage LLMs to create diverse synthetic dialogues with varying user personas to train smaller dialogue models.
+CollabLLM differs in leveraging user simulators in forward sampling to account for long-term effect in both offline and online training.
 In Table 4, we compare CollabLLM with related methods across four key dimensions.
@@
 Multiturn human-LLM collaborations are increasingly prevalent in real-world applications.
-We believe future foundation models should act as collaborators rather than passive responders, actively uncovering user intents in open-ended and complex tasks—an area where current LLMs fall short.
+Foundation models should act as collaborators rather than passive responders, actively uncovering user intents in open-ended and complex tasks—an area where current LLMs fall short.
 The key insight of CollabLLM is making LLMs more multiturn-aware by using forward sampling to estimate the long-term impact of responses.
 Through extensive simulated and real-world evaluations, we demonstrate that CollabLLM is highly effective, efficient, and engaging, while also generalizing well to new tasks and interactions, advancing the frontiers of human-centered LLMs.
+Acknowledgments
+We thank Doug Burger, Vishal Chowder, Jeevana Priya Inala, Hoifung Poon, Swadheen Shukla, Chandan Singh, Desney Tan and Chenglong Wang, as well as members of the Deep Learning and Health Futures groups at Microsoft Research for helpful discussions.
+We thank lab members in Leskovec and Zou’s labs for discussions and for providing feedback.
+We also gratefully acknowledge the support of NSF under Nos.
+OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM), IIS-2403318 (III); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Stanford Institute for Human-Centered AI, Chan Zuckerberg Initiative, Amazon, Genentech, GSK, Hitachi, SAP, and UCB.
 Impact Statement
@@
 This consideration would also help increase the adoption of AI in safety-critical scenarios, such as medical decision-making tasks (Liu et al., 2024), in which we believe AI models should be inclined to seek confirmation or verification (Gero et al., 2023) from an expert in case of uncertainty—a behavior that is mostly absent in current state-of-the-art LLMs.
-This work presents one of the first attempts to train LLMs in such human-centric environments.
-To promote future research in this societally beneficial direction, we plan to release all the code, models, data, benchmarks, and user simulators described in this work.
+Since the models in this work are trained collaboratively and aim to better align with user intent, concerns may arise regarding users with malevolent goals.
+However, we argue that CollabLLM can help mitigate safety risks in such cases—at least when used with LLMs that have been aligned for safety (as is the case for all models used in this work).
+Safety-aligned LLMs generally refuse to respond to unsafe queries, which often leads malicious users to obscure their true intentions in order to bypass safeguards.
+This is where our approach offers an advantage: CollabLLM often seeks to clarify user intent, creating additional opportunities to detect misuse.
+For example, malicious users might unintentionally reveal their actual goals, or their vagueness and refusal to disclose motivations could raise red flags—potentially providing the LLM with further cues for identifying unsafe behavior.
+As presented in Appendix C, we conducted various safety experiments and show that CollabLLM performs no worse than an equivalent non-collaboratively trained model in terms of safety.
 The data collected in our study involves human participants recruited through Mechanical Turk.
@@
 Given that conversations averaged 28.4 minutes, including break times, this means workers were paid more than $20 per hour on average—above the minimum wage in the country where the data was collected.
-Acknowledgments
-We thank Vishal Chowdhary, Hoifung Poon, Swadheen Shukla, Chandan Singh and Chenglong Wang, as well as members of the Deep Learning and Health Futures groups at Microsoft Research for helpful discussions.
-We thank lab members in Leskovec and Zou’s labs for discussions and for providing feedback.
+This work presents one of the first attempts to train LLMs in such human-centric environments.
+To promote future research in this societally beneficial direction, we release all the code, models, data, benchmarks, and user simulators described in this work.
 Appendix A Supplementary Discussion
@@
 Existing multiturn trajectory-based training frameworks (Shani et al., 2024; Zhou et al., 2024; Gao et al., 2024) primarily rely on learning from observed trajectory-level rewards.
-These methods estimate the utility of responses by assigning rewards post hoc to completed conversations, training models to imitate historically successful responses.
-However, this approach is fundamentally correlational—it captures statistical associations between responses and final outcomes but does not explicitly model how different responses causally influence future turns.
-In contrast, our Multiturn-aware Reward (MR) framework intervenes on model responses and uses forward sampling to explicitly simulate alternative future trajectories.
-This enables the model to estimate the counterfactual impact of different responses and adjust its behavior accordingly.
-By leveraging causal effect estimation, MR training moves beyond passive imitation of high-reward conversations and instead actively optimizes responses to maximize long-term goal attainment.
-This distinction is crucial in dynamic human-LLM interactions where user needs evolve throughout a conversation.
-Appendix B Dataset and Training Details
+These methods estimate the utility of responses by assigning rewards post hoc to completed conversations, typically training models to prefer higher-rated conversations over lower-rated ones.
+However, this approach is fundamentally observational—it captures statistical associations between responses and final outcomes, without disentangling how individual responses causally influence future turns.
+For example, in MTPO (Shani et al., 2024), the learning signal remains coarse-grained: rewards are assigned at the trajectory level, and the influence of specific turns within a conversation remains confounded and indirect.
+In contrast, our Multiturn-aware Reward (MR) framework intervenes on individual model responses and uses forward simulation to generate alternative future trajectories.
+This allows the model to estimate the counterfactual impact of different responses at each turn, thereby enabling fine-grained optimization.
+By leveraging causal effect estimation, MR training moves beyond passive imitation of high-reward conversations and instead actively selects responses to maximize long-term task success.
+This interventional approach provides turn-level credit assignment that is critical in dynamic human-LLM interactions, where user needs evolve and the consequences of early decisions compound over time.
+Appendix B Experimental Details
 B.1 Dataset Generation for Offline Training
@@
 B.2 Training Details
-We provide the hyperparameters for CollabLLM fine-tuning in Table 6.
+Hyperparameters (Table 6). We provide the hyperparameters for CollabLLM fine-tuning.
 Notably, CollabLLM relies on a minimal set of hyperparameters, using the same window size and sample size for computing MRs across multiple datasets.
@@
 | Penalty | 1e-4 | 5e-4 | 5e-4 |
-Appendix C Prompts
-C.1 User Simulator
-C.2 Prompt for Proactive Base
-C.3 System Prompt
-C.4 Interactivity Metric by LLM Judge
-C.5 Helpfulness Reward by LLM Judge
-Appendix D Question Template and Example on Abg-CoQA
+Training Cost (Table 7).
+We compute average statistics over 100 future conversations on MediumDocEdit-Chat, the document editing task, which incurs the highest computational overhead among the three tasks.
+The table shows that even at the largest window size (), the total per-sample cost remains low, suggesting that our multi-turn training setup is financially practical.
+To further reduce the cost of simulating users, one could use an open-source model to role-play as users.
+Unfortunately, at the current stage, we find that open-source models generally perform poorly, often getting “confused” and starting to solve problems as an assistant rather than acting as a user.
+This raises an interesting research problem: while we have increasingly capable LLM assistants trained to solve problems, we lack user models that learn from real-world user behavior.
+Building better user models could be valuable for running simulations in real-world applications.
+| 0.89 | 0.42 | 7.41 | 1.85 | 0.26 | 0.00174 | |||||||||||||
+| 2.55 | 0.91 | 15.84 | 4.55 | 0.69 | 0.00439 | |||||||||||||
+| 4.13 | 1.22 | 21.72 | 7.18 | 1.06 | 0.00685 |
+B.3 Full Ablation Results
+To further understand the source of performance improvements, we conduct a full ablation by training models with isolated reward signals—Helpfulness, Extrinsic Only, and Intrinsic Only—across window sizes .
+The resulting BLEU, token usage, and ITR scores are reported in Figure 9.
+We make three key observations:
+Helpfulness alone leads to marginal improvements in BLEU and ITR, but significantly increases token usage, especially at larger window sizes, suggesting verbosity rather than improved efficiency or interactivity.
+Extrinsic-only reward achieves strong BLEU scores (e.g., 0.377 at ), indicating good task alignment. However, it underperforms in ITR and often generates longer responses.
+Intrinsic-only reward improves ITR at (e.g., 0.74), but offers lower BLEU and comparable or slightly lower token efficiency, indicating better interactivity at the expense of task success.
+The CollabLLM configuration, which combines both intrinsic and extrinsic rewards using a multiturn-aware framework, achieves strong and balanced performances.
+Note that the choice of reward type (intrinsic or extrinsic) is independent of the multiturn-aware reward design.
+In practice, one can flexibly plug in different reward signals, which are then used to evaluate the responses’ long-term impact through forward sampling.
+Appendix C Safety Evaluation
+As the models in this work are collaboratively trained and designed to be more aligned with the user’s intent, concerns may arise if a user happens to have malevolent intentions.
+However, we note that CollabLLM models were finetuned from Llama-3.1-8B, which has been aligned for safety—so jailbreaking CollabLLM still poses a significant challenge.
+To determine whether collaborative training weakens the safety features inherent to a model (Llama-3.1-8B) that has undergone significant alignment steps for safety, we performed an adversarial evaluation using the Azure AI Evaluation SDK‡‡‡https://learn.microsoft.com/en-us/python/api/overview/azure/ai-evaluation-readme and prompted both the baseline and CollabLLM with various offensive queries intended to elicit unsafe responses.
+Specifically, we performed the following steps:
+Adversarial query selection: We used the SDK’s AdversarialSimulator to generate adversarial queries (e.g., queries encouraging the LLM to produce hateful comments).
+We then used the SDK’s harm evaluators (ViolenceEvaluator, SexualEvaluator, SelfHarmEvaluator, HateUnfairnessEvaluator) to categorize each query into one of four harm types: violence, sexual, self-harm, and hate.
+For each query, we used the highest score among the four evaluators to determine its harm category.
+We randomly selected 20 adversarial queries per harm category, resulting in a total of 80 queries.
+Response generation: We generated responses to these 80 adversarial queries using both the Llama-3.1-8B baseline model and CollabLLM.
+Harm scoring: We evaluated each model-generated response using all four harm evaluators to ensure comprehensive assessment.
+| Model | Harm score (0–7 range, ) | |||
+| Violence | Sexual | Self-harm | Hate | |
+| Llama-3.1-8B | 0.88 | 0.96 | 0.89 | 1.01 |
+| CollabLLM | 0.95 | 0.94 | 1.00 | 0.99 |
+The main safety results are shown in Table 8, which presents the average harm scores across the four categories.
+Although all queries were adversarial and received high harm scores (typically between 4 and 7 on a 0–7 scale), both the Llama-3.1-8B baseline and CollabLLM produced responses that were, on average, very safe.
+Most scores are in the 0–1 range, which corresponds to “very low” harm.
+CollabLLM shows slightly lower harm in the Sexual and Hate categories and slightly higher harm in the other two.
+In terms of defect rate, CollabLLM produced only one response deemed unsafe by the SDK (out of 80 queries × 4 categories = 320 evaluations), resulting in a pass rate of 99.7%.
+Coincidentally, this is the same pass rate as Llama-3.1-8B, which also had one failed evaluation.
+Overall, these results are encouraging.
+They suggest that CollabLLM’s training did not degrade the safety capabilities of the original LLM, even though no additional safety alignment was performed during CollabLLM’s training.
+Appendix D Prompts
+D.1 User Simulator
+D.2 Prompt for Proactive Base
+D.3 System Prompt
+D.4 Interactivity Metric by LLM Judge
+For the prompt template below, the ITR results reported in Table 1 use weights , , and , with the final score rescaled as , as all methods achieve an average ITR score above 2.5.
+Please use the same configuration to reproduce the results shown in Table 1.
+Note that the absolute values of , , and do not affect the overall conclusions.
+In our most recent codebase, we adopt , , and to eliminate the need for rescaling.
+D.5 Helpfulness Reward by LLM Judge
+Appendix E Question Template and Example on Abg-CoQA
 We use the following prompt format for the LLMs to answer the question given a story.
@@
 The label of the above question is ambiguous since the user’s query about ‘‘Where did they go when they left?’’ could mean ‘‘Where did they go when they left the zoo?’’ or ‘‘Where did the grandparents go when they left me?’’.
-Appendix E User Study
+Appendix F User Study
+F.1 User Study Platform
 We provide screenshots of the interface used for human participants to interact with the AI assistants.
@@
 All data collection is fully anonymized to ensure user privacy.
+F.2 Analysis: Divergence Between Simulated and Real Users
+While user simulators were employed exclusively during training due to the large-scale conversation demands of our Multiturn-aware Reward computation, we provide a comparative analysis to study the divergence between user simulators and real users.
+We summarize key differences and similarities in communication patterns between real and simulated users below:
+| Differences | Similarities |
+| 1) Real users tend to use shorter, fragmented sentences with grammatical errors; simulators produce more complete and polished responses.
+| 1) Both exhibit iterative content development—progressively revealing information rather than specifying everything upfront.
+|
+| 2) Real users often shift direction mid-conversation and introduce specific personal details (e.g., “eight dogs”); simulated users remain more predictable and generic.
+| 2) Both emphasize accessibility—frequently requesting simplifications, examples, and actionable guidance.
+|
+| 3) Real users express emotion more bluntly (e.g., “that’s awful”) and use informal language, abbreviations, or incomplete thoughts; simulators respond in a more neutral and formal tone.
+| 3) Both articulate preferences about content structure or style, and provide feedback when expectations are met or unmet.
+|
+Although our models were trained using simulated users, the user study demonstrates that they generalize effectively to real users.
+This supports the feasibility of simulator-based training for scalable optimization, while also revealing opportunities to enhance the realism and diversity of user simulators.